Ron and Kathy are ticket sellers at their class play, Ron handling student tickets that sell for $2 each and Kathy selling adult tickets for $5.50 each. If their total income for 30 tickets was $105.50, How many tickets did Ron sell?
Answer by checkley77(12844)

R+K=30 OR R=30-K
2R+5.50K=105.50
2(30-K)+5.50K=105.50
60-2K+5.50K=105.50
3.5K=105.50-60
3.5K=45.50
K=45.50/3.5
K=13 $5.50 TICKETS WERE SOLD BY KATHY.
30-13=17 $2.00 TICKETS WERE SOLD BY RON.
PROOF:
2*17+5.50*13=105.50
34+71.50=105.50
105.50=105.50